### v3

@@ 2.1 Optimality conditions
We only need calculate the subgradients of $s$ at 00, i.e., $i\in\mathcal{J}(x)$. If $r^{\prime}(|0^{+}|)<\infty$, define for any $x\in\mathbb{R}$,
@@
Obviously, $\hat{s}(\cdot)$ is smooth on $\mathbb{R}$ with $\hat{s}^{\prime}(0)=r^{\prime}(|0^{+}|)>0$ and $s(x)=\max\{\hat{s}(x),-\hat{s}(x)\}$. By [42, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$. On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, meaning $r^{\prime}(|0^{+}|),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$. It follows from [42, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$. Hence, $\partial s(0)=\hat{\partial}s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
@@
Overall, we have shown that (4) is true. It then follows from [42, Proposition 10.5] that (3) is true. ∎
@@
Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point satisfy the necessary optimality conditions (6) and (7) are referred as the stationary point.
@@
Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of non-zero $\{u^{k}\}$ satisfying
@@
It is easy to know that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$ since $x^{*}$ is the stationary point of (P) and (8) holds. Therefore for sufficiently large $k$,
@@
$r(|u_{i}^{k}|)=r(|u_{i}^{k}|)-r(0)=r^{\prime}(\hat{u}_{i}^{k})|u^{k}_{i}|,\text{ \ with }\hat{u}_{i}^{k}\in(0,|u_{i}^{k}|).$
The concavity (Assumption 1) of $r$ implies that $r^{\prime}(\hat{u}_{i}^{k})\geq r^{\prime}(\|u^{k}\|_{\infty})$; consequently,
@@
This, combined with (11), yields
@@
implying $\limsup\limits_{k\to\infty}\eta^{k}=\limsup\limits_{k\to\infty}\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=\|\nabla f(x^{*})\|_{\infty}$. By (7), $\limsup\limits_{k\to\infty}\eta^{k}=\|\nabla f(x^{*})\|_{\infty}<\lambda r^{\prime}(|0^{+}|)$. Therefore, $\eta^{k}<\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})$ for sufficiently large $k$, as $\|\hat{u}^{k}\|_{\infty}\to 0$. It follows from (12) that for sufficiently large $k$,
@@ 2.2 Strict Saddle Points
In this section, we define the strict saddle point of the problem (P). In smooth settings, the strict saddle can be defined as the points with negative minimum eigenvalue of the Hessian matrix, which indicates a descent direction. For our case, Theorem 4 implies a saddle point can be defined using the nonzeros.
@@
A stationary point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathbb{R}^{|\mathcal{I}^{*}|}\setminus\{0\}$. If all stationary points of $F(x)$ are either strict saddle points or local minimums, $F$ satisfies the strict saddle property.
This definition is related to the “active strict saddle point” proposed in [23], in which it is shown that the presence of negative curvature alone in nonsmooth settings does not guarantee escape from saddle points. To address this, they define the active strict saddle point for nonsmooth problems as the saddle point on the active manifold containing the stationary point on which the function is partly smooth. Our definition also belongs to this kind, as $\mathcal{M}_{\mathcal{I}^{*}}$ is an active manifold containing $x^{*}$.
To show that the strict saddle property is mathematical generic, many works show that this property holds true for a full measure of linear perturbation [23, Theorem 2.9], [44, Theorem 2]. Namely, the perturbed function has the strict saddle property for almost all $v\in\mathbb{R}^{n}$. In our case, we also have the same results by considering the following perturbed nonsmooth regularization problem for $v\in\mathbb{R}^{n}$,
@@
To show the desired result, we need the assistance of the famous Sard’s theorem. Our proof of the strict saddle property mainly follows the proof of [45, Lemma 2.7] which is proved firstly in [46].
@@
be a manifold. Let $v:\mathcal{M}(s)\rightarrow\mathbb{R}^{n}$ be a smooth map: $[v(x)]_{\mathcal{I}}=\nabla_{\mathcal{I}}F(x)$ and $[v(x)]_{\mathcal{I}^{c}}=x$ with $\mathcal{I}=\mathcal{I}(s)$. The Jacobian of $v$ at $x$ is then given by $\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)$. By Sard’s Theorem,
$\{[\nabla_{\mathcal{I}}F(x);v_{\mathcal{I}^{c}}]\mid\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}$
$\implies$ $\{v\mid v=v(x),\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}$
$\implies$ $\mathcal{V}(s):=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathcal{M}\}\text{ has measure 0 in }\mathbb{R}^{n}.$
@@
(Unstable fixed point) Let $x^{*}$ be a fixed point of a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. $x^{*}$ is called an unstable fixed point of $T$ if $DT(x^{*})$ has at least one eigenvalue of magnitude strictly larger than 1.
The Center Stable Manifold (CSM) Theorem plays a crucial role in deriving the avoidance of strict saddle points. The CSM theorem establishes a favorable property for the fixed point iteration of a $C^{1}$ local diffeomorphism map: the set of initial points around a fixed point, which converge to the fixed point with at least one eigenvalue of $DT$ larger than $1$ in magnitude, has Lebesgue measure zero.
@@
If $\nabla T(0)$ is invertible, then $T$ is locally a diffeomorphism around 0. The CSM theorem only guarantees the avoidance of strict saddle point in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points that converge to any unstable fixed point has measure zero. This requires the map has the so-called 0-property, meaning the preimage of a measure zero set under this map still has measure zero. Such result was established in [47, Theorem 2] for a global diffeomorphism map and then for a globally lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous) in [23]. Moreover, to extend this property from a particular stationary point $x^{*}$ to the set of all stationary points, this can be shown by using the fact that $\mathbb{R}^{n}$ is second countable no matter the number of stationary points are countable (see [23, Corollary 2.12] for a standard proof). We list their results below.
@@
[47, Theorem 2] Let $g$ be a $C^{1}$ mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $det(DT(x))\neq 0$ for all $x\in\mathbb{R}^{n}$. Let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ Then the set of initial points that converge to an unstable fixed point has measure zero, $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
@@
  2. (2).
    Show the fixed point iteration is locally a $C^{1}$ local diffeomorphism around a strict saddle point (e.g., a $C^{1}$ smooth map with invertible Jacobian at fixed point), so that the CSM theorem guarantees the avoidance of converging to the strict saddle point.
  3. (3a).
    Show the fixed point iteration is globally a diffeomorphism.
  4. (3b).
    (Alternative to (3a)) Show the fixed point iteration is globally a lipeomorphism.
@@ 3 Damped Iterative Reweighted $\ell_{1}$ Algorithms
where $w(x_{i})=r^{\prime}(|x_{i}|)$. To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a stationary point and a Lipeomorphism globally. However, this is not true in our case since $S$ is generally non-invertible.
This algorithm may quickly converges to a local minimizers due to the rapid growth of the weight values, especially for the case when $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$ can be quickly encountered. To address this issue, a common approach is to add a perturbation $\epsilon_{i}>0$ to each $r$ and drive $\epsilon_{i}$ to 0 during the iteration, resulting in the following approximation $F(x,\epsilon)$ of $F(x)$
@@
$S$ is not a bijection, as it maps multiple values to 00. Motivated by the technique used in [23], we propose to use the damped operator: $T=(1-\alpha)I+\alpha S$, which can be written explicitly as
@@
As for $x_{i}^{*}=0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)]<0$ by (7). Same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$. Therefore, $S_{i}^{x}(x^{*},0)=0=x_{i}^{*}$. Overall, we have shown that $x^{*}=S^{x}(x^{*},0)$ in all cases.
@@
If $x_{i}^{*}=0$, it indicates that $|\frac{1}{\beta}\nabla_{i}f(x^{*})|=|x_{i}^{*}-\frac{1}{\beta}\nabla_{i}f(x^{*})|\leq\tfrac{\lambda}{\beta}w(0,0)=\tfrac{\lambda}{\beta}r^{\prime}(|0^{+}|)$. Hence, $\nabla_{i}f(x^{*})\in[-\lambda r^{\prime}(|0^{+}|),-\lambda r^{\prime}(|0^{+}|)]$. Overall, $x^{*}$ satisfies (5) and is a stationary point of (P). By Assumption 2, it satisfies (6) and (7).
@@ 3.1 Convergence analysis
It is noteworthy that the same conclusion for Algorithm 10 can be readily deduced by setting $\epsilon$ to zero. We first make the following assumptions for the algorithm.
@@
This assumption also implies that if $r^{\prime}(|0^{+}|)=\infty$, then for any stationary point $x^{*}$, $|x_{i}^{*}|\geq(r^{\prime})^{-1}(L_{f}/\lambda),\forall i\in\mathcal{I}^{*}$ by (6). Therefore, we know $r^{\prime\prime}(|x_{i}^{*}|)>0$ is bounded above and this is also the case for $r^{\prime}(|0^{+}|)<\infty$. Overall, we can conclude
$\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$ (20)
on all stationary points in $\mathcal{L}(F^{0})$.
@@
Since the concavity of $r(\cdot)$ on $\mathbb{R}_{++}$ gives $r(x_{1})\leq r(x_{2})+r^{\prime}(x_{2})(x_{1}-x_{2})$ for any $x_{1},x_{2}\in\mathbb{R}_{++}$, we can imply that for $i=1,...,n$,
@@
where the first inequality is by the fact that $\epsilon^{k+1}\leq\epsilon^{k}$.
@@
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, it follows from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (21) holds for any $k\in\mathbb{N}$. ∎
@@
Firstly, boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. This also implies $\text{sign}(y_{i}^{k})=\text{sign}(x_{i}^{*}),i\in\mathcal{I}^{*}$ for sufficiently large $k\in\mathcal{S}$. The optimality condition of the $k$th subproblem implies

### v4

@@ 2.1 Optimality conditions
We only need to calculate the subgradients of $s$ at 00, i.e., $i\in\mathcal{J}(x)$. If $r^{\prime}(|0^{+}|)<\infty$, define for any $x\in\mathbb{R}$,
@@
Obviously, $\hat{s}(\cdot)$ is smooth on $\mathbb{R}$ with $\hat{s}^{\prime}(0)=r^{\prime}(|0^{+}|)>0$ and $s(x)=\max\{\hat{s}(x),-\hat{s}(x)\}$. By [42, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$. On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, which means $r^{\prime}(|0^{+}|)\in\hat{\partial}s(0),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$. It follows from [42, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$. Hence, $\partial s(0)=\hat{\partial}s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
@@
In general, we have shown that (4) is true. It then follows from [42, Proposition 10.5] that (3) is true. ∎
@@
Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point that satisfies the necessary optimality conditions (6) and (7) is referred to as the stationary point.
@@
Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of nonzero $\{u^{k}\}$ satisfying
@@
From the classic optimality condition for smooth function and (9), it is easy to see that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$. Therefore for sufficiently large $k$,
@@
$r(|u_{i}^{k}|)=r(|u_{i}^{k}|)-r(0)=r^{\prime}(\hat{u}_{i}^{k})|u^{k}_{i}|,\text{ \ with }\hat{u}_{i}^{k}\in(0,|u_{i}^{k}|).$ (13)
The concavity (Assumption 1) of $r$ implies that $r^{\prime}(\hat{u}_{i}^{k})\geq r^{\prime}(\|u^{k}\|_{\infty})$, which, together with (13), yields that
@@
This, combined with (12) and (11), yields
@@
implying $\limsup\limits_{k\to\infty}\eta^{k}=\limsup\limits_{k\to\infty}\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}\leq\|\nabla f(x^{*})\|_{\infty}$. By (7), $\limsup\limits_{k\to\infty}\eta^{k}\leq\|\nabla f(x^{*})\|_{\infty}<\lambda r^{\prime}(|0^{+}|)$. Therefore, $\eta^{k}<\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})$ for sufficiently large $k$, as $\|\hat{u}^{k}\|_{\infty}\to 0$. It follows from (14) that for sufficiently large $k$,
@@ 2.2 Strict Saddle Points
In this section, we define the strict saddle point of the problem (P). In smooth settings, the strict saddle can be defined as the points with negative minimum eigenvalue of the Hessian matrix, which indicates a descent direction. For our case, Theorem 4 implies that a saddle point can be defined using nonzeros.
@@
A stationary point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathbb{R}^{|\mathcal{I}^{*}|}\setminus\{0\}$. If all stationary points of $F(x)$ are strict saddle points or local minimums, $F$ satisfies the strict saddle property.
This definition is related to the “active strict saddle point” proposed in [23], in which it is shown that the presence of negative curvature alone in nonsmooth settings does not guarantee escape from saddle points. To address this, they define the active strict saddle point for nonsmooth problems as the saddle point on the active manifold containing the stationary point on which the function is partly smooth. Our definition also belongs to this kind, since $\mathcal{M}_{\mathcal{I}^{*}}$ is an active manifold that contains $x^{*}$.
To show that the strict saddle property is mathematically generic, many works show that this property holds for a full measure of linear perturbation [23, Theorem 2.9], [44, Theorem 2]. Namely, the perturbed function has the strict saddle property for almost all $v\in\mathbb{R}^{n}$. In our case, we also have the same results by considering the following perturbed nonsmooth regularization problem for $v\in\mathbb{R}^{n}$,
@@
To show the desired result, we need the assistance of the famous Sard’s theorem. Our proof of the strict saddle property mainly follows the proof of [45, Lemma 2.7], which is first proved in [46].
@@
be a manifold. Let $v:\mathcal{M}(s)\rightarrow\mathbb{R}^{n}$ be a smooth map: $[v(x)]_{\mathcal{I}}=\nabla_{\mathcal{I}}F(x)$ and $[v(x)]_{\mathcal{I}^{c}}=x$ with $\mathcal{I}=\mathcal{I}(s)$. The Jacobian of $v$ at $x$ is then given by $Dv(x)=\begin{bmatrix}\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)&0\\ 0&I\end{bmatrix}$, and $Dv(x)$ has full rank if and only if $\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)$ has full rank. By Sard’s Theorem,
$\{v(x)\mid\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n},$
which is equivalent to
$\mathcal{V}(s):=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}(s)|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}.$
@@
(Unstable fixed point) Let $x^{*}$ be a fixed point on a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. $x^{*}$ is called an unstable fixed point of $T$ if $DT(x^{*})$ has at least one eigenvalue of magnitude strictly larger than 1.
The Center Stable Manifold (CSM) Theorem plays a crucial role in deriving the avoidance of strict saddle points. The CSM theorem establishes a favorable property for the fixed-point iteration of a $C^{1}$ local diffeomorphism map: the set of initial points around a fixed point, which converge to the fixed point with at least one eigenvalue of $DT$ larger than $1$ in magnitude, has Lebesgue measure zero.
@@
If $\nabla T(0)$ is invertible, then $T$ is locally a diffeomorphism around 0. The CSM theorem only guarantees the avoidance of strict saddle point in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points that converge to any unstable fixed point has measure zero. This requires that the map has the so-called 0-property, meaning the preimage of a measure zero set under this map still has measure zero. This result was established in [47, Theorem 2] for a global diffeomorphism map and then for a global lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous) in [23]. Moreover, to extend this property from a particular stationary point $x^{*}$ to the set of all stationary points, this can be shown using the fact that $\mathbb{R}^{n}$ is second countable no matter if the number of stationary points is countable (see [23, Corollary 2.12] for a standard proof). We list their results below.
@@
[47, Theorem 2] Let $g$ be a $C^{1}$ mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $det(DT(x))\neq 0$ for all $x\in\mathbb{R}^{n}$. Let $\mathcal{U}_{T}$ consist of all unstable fixed points $x$ of $T$ Then the set of initial points that converge to an unstable fixed point has measure zero, $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
@@
  2. (2)
    Show that the fixed point iteration is locally a $C^{1}$ local diffeomorphism around a strict saddle point (e.g., a $C^{1}$ smooth map with invertible Jacobian at the fixed point), so that the CSM theorem guarantees the avoidance of converging to the strict saddle point.
  3. (3a)
    Show that the fixed-point iteration is globally a diffeomorphism.
  4. (3b)
    (Alternative to (3a)) Show that the fixed-point iteration is globally a lipeomorphism.
@@ 3 Damped Iterative Reweighted $\ell_{1}$ Algorithms
where $w(x_{i})=r^{\prime}(|x_{i}|)$.
This algorithm may quickly converge to local minimizers due to the rapid growth of the weight values, especially for the case where $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$ can be quickly encountered. To address this issue, a common approach is to add a perturbation $\epsilon_{i}>0$ to each $r$ and drive $\epsilon_{i}$ to 0 during the iteration, resulting in the following approximation $F(x,\epsilon)$ of $F(x)$
@@
To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a stationary point and a Lipeomorphism globally. However, this is not true in our case since $S$ is generally noninvertible. In particular, $S$ is not a bijection, as it maps multiple values to 00. Motivated by the technique used in [23], we propose to use the damped operator: $T=(1-\alpha)I+\alpha S$, which can be written explicitly as
@@
As for $x_{i}^{*}=0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)]<0$ by (7). Same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$. Therefore, $S_{i}^{x}(x^{*},0)=0=x_{i}^{*}$. In general, we have shown that $x^{*}=S^{x}(x^{*},0)$ in all cases.
@@
If $x_{i}^{*}=0$, it indicates that $|\frac{1}{\beta}\nabla_{i}f(x^{*})|=|x_{i}^{*}-\frac{1}{\beta}\nabla_{i}f(x^{*})|\leq\tfrac{\lambda}{\beta}w(0,0)=\tfrac{\lambda}{\beta}r^{\prime}(|0^{+}|)$. Hence, $\nabla_{i}f(x^{*})\in[-\lambda r^{\prime}(|0^{+}|),-\lambda r^{\prime}(|0^{+}|)]$. In general, $x^{*}$ satisfies (5) and is a stationary point of (P). By Assumption 2, it satisfies (6) and (7).
@@ 3.1 Convergence analysis
It should be noted that the same conclusion for Algorithm 10 can be readily deduced by setting $\epsilon$ to zero. We first make the following assumptions for the algorithm.
@@
This assumption also implies that if $r^{\prime}(|0^{+}|)=\infty$, then for any stationary point $x^{*}$, $|x_{i}^{*}|\geq(r^{\prime})^{-1}(L_{f}/\lambda),\forall i\in\mathcal{I}^{*}$ by (6). Therefore, we know $r^{\prime\prime}(|x_{i}^{*}|)>0$ is bounded above and this is also the case for $r^{\prime}(|0^{+}|)<\infty$. Overall, we can conclude that for any stationary point $x^{*}\in\mathcal{L}(F^{0})$, there exists $\rho$ such that
$\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho.$ (22)
@@
Suppose Assumptions 1-3 hold. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 10 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{f}}{2}$. It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and that the reduction satisfies
@@
Since the concavity of $r(\cdot)$ on $\mathbb{R}_{++}$ gives $r(x_{1})\leq r(x_{2})+r^{\prime}(x_{2})(x_{1}-x_{2})$ for any $x_{1},x_{2}\in\mathbb{R}_{++}$, we have that for $i=1,...,n$,
@@
where the first inequality is due to the fact that $\epsilon^{k+1}\leq\epsilon^{k}$.
@@
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (23) holds for any $k\in\mathbb{N}$. ∎
@@
Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. This also implies $\text{sign}(y_{i}^{k})=\text{sign}(x_{i}^{*}),i\in\mathcal{I}^{*}$ for sufficiently large $k\in\mathcal{S}$. The optimality condition of the $k$th subproblem implies